Measurement of the left- and right- hand gripping strengths of 9 lefthanded writers are recorded:

Person Left hand Right hand || 1 2 | 140 89 138 87 3 4 5 6 7 8 125 134 146 106 117 86 113 126 130 98 110 83 9 79 69\We want to test whether the left- hand gripping strength and the right hand gripping strengths of left-handed writers are different at 1% level of significance. Instead of performing the test, construct a confidence interval to draw a conclusion about the test.
Answer #1

solution Here Ho Md=0 vs Hild to from the given data n=9 d = 7.56, sd = U-75, d-f=9-1=8 At 99% confidence tale - 3.355 The 99